1- find the common ratio and seventh term of the following sequence: 2÷9 ,2÷3,2,6,18....
2-the first term of a geometric series is 1 and the fourth term is 27.find the sum of the first terms.
1
Expert's answer
2019-11-26T14:00:58-0500

1.

"r=\frac{a_n}{a_{n-1}}=\frac{a_4}{a_3}=\frac{6}{2}=3.\\
a_1=\frac{2}{9},\\
a_7=a_1r^{7-1}=\frac{2}{9}*3^6=162."


2.

"a_1=1,\;\;a_4=a_1r^3=27,\\
r=3.\\
S_n=a_a\frac{1-r^n}{1-r}=\frac{1-3^n}{1-3}=\frac{3^n-1}{2}.\\
S_4=\frac{3^4-1}{2}=40."